Find the area of the region enclosed by the graphs of y= 3/x and y=4-x
1
Expert's answer
2021-11-15T07:35:48-0500

Given: "y=3/x ,\ y=4-x"

Point of intersection of these two curve will occur at

"3/x=4-x\Rightarrow x^2-4x+3=0
\\\Rightarrow (x-3)(x-1)=0
\\\Rightarrow x=1,3"

Graph of the above function is shown below




Area of the region enclosed by these two curves

"A=\int_{1}^{3} (4-x-\frac{3}{x})dx=[4x-\frac{x^2}{2}-3lnx]_{1}^{3}\\
=12-\frac{9}{2}-3ln3-4+\frac{1}{2}+3ln1=4-3ln3" ["\because ln 1=0" ]